OpenStudy (anonymous):

hi im having trouble understanding variables and solving equation. i need to simplify this -4y(3y)-5+2(5y+7) can anyone help

OpenStudy (anonymous):

You need to distribute through to simplify. What do you not understand?

OpenStudy (anonymous):

all of it im new at this i get confused on were to stsrt

OpenStudy (anonymous):

Ok. Well when you multiply variables the format is as follows. Let * be multiplication. \[x*x=x^2\] and \[x*x*x=x^3\] and so on.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

-12y²-5+10y+14

OpenStudy (anonymous):

So maybe it's easier to work through the problem, or just spit out the answer like what has been done already.

OpenStudy (anonymous):

cyter how did you come up with that . if that is the answer fine but that dose not tell me how to do it

OpenStudy (anonymous):

yes i would like to work through it step by step please so i can under stand it better

OpenStudy (anonymous):

Ok so the first part you have -4y(3y). The parenthesis just separate quantities and mean to multiple. So -4 times 3 give -12 and y times y gives y². Thus you get -12y² for the first part. Does this make sense?

OpenStudy (anonymous):

Multiply* not multiple.

OpenStudy (anonymous):

sorry thats wrong -4(3y)-5+2(5y+7) this is the correct one

OpenStudy (anonymous):

Ok well you do the same process. So -4(3y) gives -12y. All I did was multiple what is in parenthesis by -4. Then you subtract 5 giving you -12y-5+2(5y+7). To simplify the 2(5y+7) you multiply what's in parenthesis by what is right outside the parenthesis. So 2 times (5y+7). So what does 2(5y+7) equal when you multiply it out?

OpenStudy (anonymous):

10+7=17y

OpenStudy (anonymous):

Not quite. 2(5y+7)=10y+14. You multiply each thing in the parenthesis by 2.

OpenStudy (anonymous):

2*5=10 &2*7=14

OpenStudy (anonymous):

So once we have simplified all parts we get: -12y-5+10y+14. The last step is to add things that look alike.

OpenStudy (anonymous):

You are correct above. But you can't forget the y. The y is a variable and the 5 is a coefficient. When you multiply variables by a number you just get that many variables. So 3 times y gives 3y, 7 times x gives 7x.

OpenStudy (anonymous):

-12y+10y +14==12y ?

OpenStudy (anonymous):

Ok so not sure if you were understanding how we got to -12y-5+10y+14 but if we do continue when I say add like terms I mean add any regular numbers so like -5 and 14 to get 9 and add any variables that look the same. -12y and 10y look the same. Both have a y and both have coefficients. So -12y+10y=-2y I added the -12 and 10 and just kept the y. Final answer should be -2y+9

OpenStudy (anonymous):

ok can we try another one

OpenStudy (anonymous):

Sure.

OpenStudy (anonymous):

-10+4(-3b+3)+2(6b-1)

OpenStudy (anonymous):

Now remember what I said. The number in front of the parenthesis is multiplied to each term inside the parenthesis and if there is a variable, b for example, do not leave it off.

OpenStudy (anonymous):

okdo i start with 4*-3 or 2*6b or dose it matter

OpenStudy (anonymous):

You can start wherever you like. I normally work from the left to the right so I don't forget anything :)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

4*(-3b)=12b

OpenStudy (anonymous):

Close. Don't forget the negative sign. So -12b

OpenStudy (anonymous):

4*3=12

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

(-12b)+12+12b-11

OpenStudy (anonymous):

should it look like that

OpenStudy (anonymous):

or it should be a -2 at the end

OpenStudy (anonymous):

Should be a -2 at the end and don't forget the -10 at the beginning. So after simplifying a little you get -10-12b+12+12b-2

OpenStudy (anonymous):

now we add the like #first

OpenStudy (anonymous):

Yes now you add the like terms.

OpenStudy (anonymous):

-12b+12b=-24b

OpenStudy (anonymous):

You have -12b+12b so you have a negative 12b plus a positive 12b what does that give you? Not -24b.

OpenStudy (anonymous):

+24b

OpenStudy (anonymous):

Nope. If you have 12 of something and then take away 12 of something you are left with zero. -12b+12b=0

OpenStudy (anonymous):

-10+24b+12

OpenStudy (anonymous):

-10-12b+12+12b-2=-10-12b+12b+10=0

OpenStudy (anonymous):

rewrite ok so the answer is 0

OpenStudy (anonymous):

Yep :)

OpenStudy (anonymous):

one more ok

OpenStudy (anonymous):

Ok.

OpenStudy (anonymous):

-5(-x+2)+8(-x)+3(-2x-2)+16

OpenStudy (anonymous):

-5*2=-10

OpenStudy (anonymous):

no no oh -5*1=-4

OpenStudy (anonymous):

-5*2=-10

OpenStudy (anonymous):

-5 x -1 = 5

OpenStudy (anonymous):

i thought that was how it went wrong rule i gess

OpenStudy (anonymous):

Uh maybe you meant -5+1=-4 that's correct :)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

then-5*2=10

OpenStudy (anonymous):

-5*2=-10 A negative number times a positive number is a negative number. A negative number times a negative number is a positive number.

OpenStudy (anonymous):

ok -10

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

so-4+-10=-14

OpenStudy (anonymous):

should it read -14+-7+6x+6+16 i hope!!

OpenStudy (anonymous):

i for got the perentacies around the -7 because of the two signs together

OpenStudy (anonymous):

I get 5x-10-8x-6x-6

OpenStudy (anonymous):

Forgot the +16 on the end. So 5x-10x-8x-6x-6+16

OpenStudy (anonymous):

so were did i go wrong

OpenStudy (anonymous):

doesn't (-x) mean 1

OpenStudy (anonymous):

or -1

OpenStudy (anonymous):

Mean 1, no. There is a 1 coefficient in front of the x. So -x is the same as -1x.

OpenStudy (anonymous):

so wouldn't that be 8*-1x=-7

OpenStudy (anonymous):

-7x oops

OpenStudy (anonymous):

No. You are multiplying, not adding. 8*(-1x)=-8x if you cannot add these terms since they aren't "like" terms.

OpenStudy (anonymous):

I'm going to go. But I will leave you with the work for the problem and maybe from what we've discussed you can figure it out.

OpenStudy (anonymous):

ok i told you im confused haha

OpenStudy (anonymous):

i hope so thanks for your help

OpenStudy (anonymous):

-5(-x+2)+8(-x)+3(-2x-2)+16 5x-10-8x-6x-6+16 here I just multiplied through -9x+0 here I added like terms (added the numbers with the x's by themselves to get -9x and then the numbers by themselves to get -10-6+16=0